# FC-KAN: Function Combinations in Kolmogorov-Arnold Networks

### v2

@@ 3.4 FC-KAN
which $f_{DoG}$ and $f_{BS}$ refer to DoG and B-spline functions. Finally, we use the combined output to compute the cross-entropy loss against the true labels and train the models.
@@ 4.1 Training Configuration
There are 5 independent training runs for each model on the MNIST [60] and Fashion-MNIST [61] datasets to obtain a more reliable overall performance assessment. We then select the best metric value from all runs to minimize the impact of training variability and accurately gauge the models’ maximum potential. To maintain simplicity in the network design, we utilized only activation functions (SiLU), linear transformations, and layer normalization in all models: BSRBF-KAN, EfficientKAN, FastKAN, FasterKAN, FC-KAN, and MLP. We do not use LiuKAN because its design, as the author intended, results in longer training times [1].
Table 1: The number of parameters by models.
@@
In FC-KAN models, we combine 2 out of 4 functions: B-splines (denoted as BS), Radial Basis Functions (denoted as RBF, specifically using GRBFs), Difference of Gaussians (denoted as DoG), and base linear transformations (denoted as BASE), to create 6 FC-KAN variants. All variants use a quadratic function representation in the output for the experiments. The FC-KAN models are: FC-KAN (DoG+BS), FC-KAN (DoG+RBF), FC-KAN (DoG+BASE), FC-KAN (BS+RBF), FC-KAN (BS+BASE), and FC-KAN (RBF+BASE).
@@ 4.2 Model Performance
Figure 4 shows the training losses, represented on a logarithmic scale, for MLP and KAN models on the MNIST and Fashion-MNIST datasets. The loss performance of each model was evaluated based on an independent training run. FC-KAN consistently achieves the lowest training losses across both datasets, followed by BSRBF-KAN due to its fast convergence feature. In contrast, Faster-KAN records the highest training loss on MNIST, while MLP performs similarly on Fashion-MNIST.
@@
FC-KAN (DoG+RBF) 100.00 ± 0.00 97.83 ± 0.04 97.81 ± 0.04 244.26
FC-KAN (DoG+BASE) 99.94 ± 0.04 97.81 ± 0.03 97.78 ± 0.03 234.13
FC-KAN (BS+RBF) 99.95 ± 0.03 97.74 ± 0.04 97.71 ± 0.04 234.07
FC-KAN (BS+BASE) 99.96 ± 0.02 97.78 ± 0.03 97.75 ± 0.03 235.04
FC-KAN (RBF+BASE) 99.97 ± 0.02 97.79 ± 0.03 97.76 ± 0.03 228.16
@@
FC-KAN (DoG+RBF) 99.68 ± 0.08 89.93 ± 0.07 89.87 ± 0.07 339.51
FC-KAN (DoG+BASE) 98.24 ± 0.55 89.81 ± 0.07 89.74 ± 0.07 326.38
FC-KAN (BS+RBF) 98.58 ± 0.43 89.72 ± 0.07 89.67 ± 0.06 327.49
FC-KAN (BS+BASE) 98.81 ± 0.35 89.75 ± 0.06 89.70 ± 0.06 327.31
FC-KAN (RBF+BASE) 98.97 ± 0.30 89.74 ± 0.05 89.69 ± 0.05 318.95
@@
In general, FC-KAN models outperformed others on MNIST and Fashion-MNIST but require more training time due to the quadratic function representation for output combination, as shown in Table 2. This trade-off between training time and model performance is considered reasonable. The best-performing model is FC-KAN (DoG+BS), which achieved a validation accuracy of 97.91% on MNIST and 89.99% on Fashion-MNIST. Although it has the lowest performance on Fashion-MNIST, the MLP model has the fastest training time and shows competitive performance on MNIST, even outperforming BSRBF-KAN, FastKAN, FasterKAN, and EfficientKAN on this dataset.
BSRBF-KAN, FC-KAN (DoG+BS), and FC-KAN (DoG+RBF) exhibit the best convergence on MNIST, while FC-KAN (DoG+RBF) performs the best on Fashion-MNIST, followed by FC-KAN (DoG+BS) and BSRBF-KAN. We observe that fast convergence is achieved in KAN models that incorporate function combinations rather than relying on single functions. This finding is important to consider when designing KANs with a focus on achieving rapid convergence.
@@ 4.3 Comparison of Combination Methods
From the results in Table 3, the quadratic function best represents the combination output and outperforms other combinations, although its models require more training time. It is clear that the outputs combined by element-wise operations consistently outperform other methods, demonstrating superior accuracy. This indicates that element-wise combinations are more effective in capturing and integrating relevant features from the data, leading to better performance. Meanwhile, the output combination by concatenation shows the worst results. The linear transformation applied to the concatenated output outperforms concatenation alone, but still achieves only average performance. Similarly, the maximum, minimum, and average outputs do not deliver superior results.
@@ 4.4 Model Performance with Limited Data
Figure 5 illustrates the performance of the FC-KAN models across different training subsets and function combinations. In the MNIST dataset, the RBF+BASE combination demonstrates superior performance with 1%, 5%, and 10% of the data; however, the highest performance on the full training data is recorded for the DoG+BS combination. This observation underscores the challenges of predicting the optimal function combination when working with limited training data. In the Fashion-MNIST dataset, DoG+BS achieves the best performance with 1% of the data, while DoG+RBF excels with both 5% and 10%. These results suggest that DoG+BS and DoG+RBF may perform well with the full training dataset. Indeed, DoG+BS consistently outperforms other combinations in full training, with DoG+RBF serving as a strong competitor, as in Table 2. Overall, our findings indicate that accurately predicting the performance of FC variants on the complete training dataset based solely on their performance with smaller subsets is challenging. This variability may depend on the specific datasets and the portions of data used for training.
@@ 6 Conclusion
We introduced FC-KAN, which uses various popular mathematical functions to represent data features and combines their outputs using different methods, primarily through element-wise operations in low-dimensional layers, to address image classification problems. In the experiments, we designed FC-KAN to combine pairs of functions, such as B-splines, wavelets, and radial basis functions, using several output combinations on the MNIST and Fashion-MNIST datasets. We found that FC-KAN outperformed MLP and other KAN models with the same network structure, although it required more training time. Among the variants, FC-KAN (DoG+BS), which combines the Difference of Gaussians and B-splines with a quadratic function representation in the output, achieved the best results on both datasets. Our experiments also show that it is not easy to detect the function combination performance in the full training data based on small-scale experiments. FC-KAN demonstrates promising potential in using function combinations to design KANs and improve model performance. However, there are several aspects, such as exploring other functions and combinations to extract data features for further model enhancement, that we would like to investigate. These will be the focus of our future work.

### v3

@@ 3.4 FC-KAN
$=\mathbf{o}_{DoG}+\mathbf{o}_{BS}+\mathbf{o}_{DoG}\odot\mathbf{o}_{BS}+\mathbf{o}_{DoG}\odot\mathbf{o}_{DoG}+\mathbf{o}_{BS}\odot\mathbf{o}_{BS}$
which $f_{DoG}$ and $f_{BS}$ refer to DoG and B-spline functions. Finally, we use the combined output to compute the cross-entropy loss against the true labels when training the models.
@@ 4.1 Training Configuration
There are 5 independent training runs for each model on the MNIST [62] and Fashion-MNIST [63] datasets to obtain a more reliable overall performance assessment. We then calculate the average value from all runs to minimize the impact of training variability and accurately gauge the models’ maximum potential. To maintain simplicity in the network design, we utilized only activation functions (SiLU), linear transformations, and layer normalization in all models: BSRBF-KAN, EfficientKAN, FastKAN, FasterKAN, FC-KAN, and MLP. We do not use LiuKAN because its design, as the author intended, results in longer training times [1].
Table 1: The number of parameters by models. The number of parameters includes both used and unused parameters.
@@
In FC-KAN models, we combine 2 out of 4 functions: B-splines (denoted as BS), Radial Basis Functions (denoted as RBF, specifically using GRBFs), Difference of Gaussians (denoted as DoG), and linear transformations (denoted as BASE), to create 6 FC-KAN variants. All variants use a quadratic function representation in the output for the experiments. The FC-KAN models are: FC-KAN (DoG+BS), FC-KAN (DoG+RBF), FC-KAN (DoG+BASE), FC-KAN (BS+RBF), FC-KAN (BS+BASE), and FC-KAN (RBF+BASE).
@@ 4.2 Model Performance
FC-KAN (DoG+RBF) 100.00 ± 0.00 97.76 ± 0.04 97.73 ± 0.04 225.23
FC-KAN (DoG+BASE) 99.82 ± 0.11 97.76 ± 0.02 97.73 ± 0.02 213.87
FC-KAN (BS+RBF) $9$ 0.00 97.53 ± 0.04 97.49 ± 0.04 233.89
FC-KAN (BS+BASE) 100.00 ± 0.00 97.93 ± 0.05 97.91 ± 0.05 238.94
FC-KAN (RBF+BASE) 100.00 ± 0.00 97.85 ± 0.03 97.82 ± 0.04 193.71
@@
FC-KAN (DoG+RBF) 99.82 ± 0.03 89.86 ± 0.12 89.81 ± 0.12 309.81
FC-KAN (DoG+BASE) 95.36 ± 0.13 89.57 ± 0.07 89.49 ± 0.07 300.13
FC-KAN (BS+RBF) 99.60 ± 0.09 89.45 ± 0.10 89.43 ± 0.10 330.82
FC-KAN (BS+BASE) 99.73 ± 0.02 89.90 ± 0.09 89.85 ± 0.10 326.57
FC-KAN (RBF+BASE) 99.79 ± 0.03 89.69 ± 0.03 89.65 ± 0.04 277.13
@@
Figure 5 shows the training losses, represented on a logarithmic scale, for MLP and KAN models on the MNIST and Fashion-MNIST datasets. The loss performance of each model was evaluated based on an independent training run. FC-KAN consistently achieves the lowest training losses across both datasets, followed by BSRBF-KAN due to its fast convergence feature. In contrast, FasterKAN records the highest training loss on MNIST, while MLP performs similarly on Fashion-MNIST.
In general, FC-KAN models outperformed others on MNIST and Fashion-MNIST but require more training time due to the quadratic function representation for output combination, as shown in Section 4.2. This trade-off between training time and model performance is considered reasonable. The best-performing models are FC-KAN (BS+BASE) and FC-KAN (DoG+BS), which achieved validation accuracies of 97.93% on MNIST and 89.99% on Fashion-MNIST, respectively. When calculating the metric values for both datasets, FC-KAN (DoG+BS) slightly surpassed FC-KAN (BS+BASE) and outperformed the other models. However, FC-KAN (BS+BASE) models take between 9.25% and 11.55% less training time.
Although it has the lowest performance on Fashion-MNIST, the MLP model has the fastest training time and demonstrates competitive accuracy on MNIST, even outperforming BSRBF-KAN, FastKAN, FasterKAN, and EfficientKAN on this dataset. On MNIST, MLP also contributes to the success of FC-KAN (BS+BASE), which combines outputs of B-splines and linear transformations.
BSRBF-KAN, FC-KAN (DoG+BS), FC-KAN (DoG+RBF), FC-KAN (BS+BASE), and FC-KAN (RBF+BASE) exhibit the best convergence on MNIST, while FC-KAN (DoG+RBF) performs the best on Fashion-MNIST, followed by FC-KAN (RBF+BASE) and FC-KAN (BS+BASE). We observe that fast convergence is achieved in KAN models that incorporate function combinations rather than relying on single functions. This finding is important to consider when designing KANs with a focus on achieving rapid convergence.
@@ 4.3 Comparison of Combination Methods
From the results in Section 4.3, the quadratic function best represents the combination output and outperforms other combinations, although its models require more training time. It is clear that the outputs combined by element-wise operations consistently outperform other methods, demonstrating superior accuracy. This indicates that element-wise combinations are more effective in capturing and integrating relevant features from the data, leading to better performance. Meanwhile, the output combination by concatenation shows the worst results. The linear transformation applied to the concatenated output outperforms concatenation alone, but still achieves only average performance. Similarly, the maximum, minimum, and average outputs do not deliver superior results.
@@
4.4 Misclassification Analysis
Figure 6: Heatmaps of the misclassified images in the validation set by models over MNIST and Fashion-MNIST.
To evaluate model performance across classes, we conducted a qualitative analysis of misclassifications on the validation set of the MNIST and Fashion-MNIST datasets. We selected FC-KAN (DoG+BS) along with other KANs for comparison, with each model trained for only 1 run. For each model, we counted the raw frequency of misclassified images per output class. Both MNIST and Fashion-MNIST have 10 output classes. In MNIST, the classes range from 0 to 9, while in Fashion-MNIST, they include "T-shirt/top", "Trouser", "Pullover", "Dress", "Coat", "Sandal", "Shirt", "Sneaker", "Bag", and "Ankle boot".
In MNIST, FC-KAN generally exhibits the fewest misclassification errors by class, while other models show their own weaknesses. For example, MLP and EfficientKAN performed poorly on Class 5, FastKAN on Classes 3 and 4, and BSRBF-KAN on Class 0. Class 1 had the fewest errors, while all models struggled with certain classes, such as Classes 3, 7, and 9. It is surprising in the case of Class 7, as its images seemed easy to recognize from our perspective.
In Fashion-MNIST, the performance of models by class is generally similar, but FC-KAN still outperforms in some classes, such as "T-shirt/top", "Pullover", and "Shirt." Models perform very well in certain classes, such as "Trouser", "Sandal", "Sneaker", "Bag", and "Ankle boot", while they struggle more with recognizing images belonging to "T-shirt/top", "Pullover", "Coat", and "Shirt" due to their similar appearance. This can be referred to as classification ambiguity due to the nature of the data.
In short, FC-KAN outperforms other KANs, but it only mitigates, not completely resolves, the misclassification errors found in the other models. This analysis is also helpful for focusing on design methods to address the most challenging classes and improve recognition accuracy.
@@ 4.5 Model Performance with Limited Data
Figure 7 illustrates the performance of the FC-KAN models across different training subsets and function combinations. In the MNIST dataset, the RBF+BASE combination demonstrates superior performance with 1%, 5%, and 10% of the data; however, the highest performance on the full training data is recorded for the BS+BASE combination. This observation underscores the challenges of predicting the optimal function combination when working with limited training data. In the Fashion-MNIST dataset, DoG+BS achieves the best performance with 1% of the data, while DoG+RBF excels with both 5% and 10%. These results suggest that DoG+BS and DoG+RBF may perform well with the full training dataset. Indeed, DoG+BS consistently outperforms other combinations in full training, with DoG+RBF serving as a strong competitor, as in Section 4.2. Overall, our findings indicate that accurately predicting the performance of FC variants on the complete training dataset based solely on their performance with smaller subsets is challenging. This variability may depend on the specific datasets and the portions of data used for training.
@@ 6 Conclusion
We introduced FC-KAN, which uses various popular mathematical functions to represent data features and combines their outputs using different methods, primarily through element-wise operations in low-dimensional layers, to address image classification problems. In the experiments, we designed FC-KAN to combine pairs of functions, such as B-splines, wavelets, and radial basis functions, using several output combinations on the MNIST and Fashion-MNIST datasets.
We found that FC-KAN outperformed MLP and other KAN models in terms of accuracy using the same network structure, although it required more training time. This is supported by a misclassification analysis, where FC-KAN achieves the fewest errors per class but still exhibits classification ambiguity errors, similar to other models. Among the variants, FC-KAN (DoG+BS) and FC-KAN (BS+BASE), which combine DoGs and B-splines, as well as MLPs and B-splines, respectively, in a quadratic function representation of the output, achieved the best results on both datasets. Our experiments also show that it is not easy to detect the function combination performance in the full training data based on small-scale experiments.
FC-KAN has promising potential for utilizing function combinations to design KANs and enhance model performance. However, we will aim to investigate several aspects further. These include exploring alternative functions and combinations for feature extraction, developing methods to reduce parameter usage while maintaining or improving model performance, and examining the impact of different function combinations on the stability and efficiency of KANs. These will be the focus of our future work.